# On the singular position-dependent mass

### v1

@@ I Introduction
where $F(r,t)$ is the envelope function, $E(k)$ the dispersion energy, $k$ the crystal momentum, and $V(r)$ the potential due to external sources (impurities). Nonetheless, one highlights that considering a many-body Hamiltonian approximation to an electron, it is possible to derive a Schrödinger-like equation with PDM. In this context, the PDM concept has been the subject of intensive study in several systems with effective mass Mustafa , Costa , Zare , Ghafourian , Pourali . Furthermore, it is noteworthy that significant recent technological and theoretical advances have motivated considerable investigation into systems with PDM. Such interest arises from the crucial role played by this property, which enables the analysis of semiconductor impurities Luttinger , Slater , Wannier , nonlinear optical phenomena Guo , Zhang , and quantum-relativistic problems pacheco , ILima0 .
@@ II The Generalized Schrödinger theory
Therefore, to avoid ambiguity arising from the symmetrization of the kinetic energy operator, allow us to consider Schrödinger’s equation exposed in Eq. (7). So, by analyzing Eq. (7) with the results presented in the literature, we conclude that our theory is equivalent to the ordered BenDaniel-Duke PDM problem BenDaniel . The BenDaniel-Duke theory has proven to be suitable for studying nanostructures subject to impurities, displacements, and geometric imperfections Willatzen .
@@ III The singular PDM
In this article, we consider a singular mass profile. Singular mass arises in cosmology when studying astrophysical objects such as black holes Maluf . In our study, we are interested in exploring the physical implications of this system within a non-relativistic one-dimensional theory. Thus, let us consider a one-dimensional crystalline lattice as illustrated in Fig. 1.
@@
Here, $m_{0}$ has mass dimensions and describes the intensity of the effective mass. Furthermore, $x_{0}$ represents the equilibrium spacing between lattice sites. We present the PDM varying $m_{0}$ (keeping $x_{0}$ constant) and varying $x_{0}$ (keeping $m_{0}$ constant), respectively, in Fig. 2[(a) and (b)].
@@ III.1 Connection with solid-state physics
Allow us to attempt to establish a phenomenological connection with our model. To accomplish our purpose, let us recall that the PDM has an equivalent distribution in reciprocal space. Thereby, the reciprocal mass is
@@ III.2 The eigenstates
Naturally, one assumes that the atoms in the crystalline lattice suffer the action of a restoring force Gomes , Kittel , i.e., the atoms belonging to the lattice behave as a set of harmonic oscillators. Therefore, we adopt an interaction harmonic to the description of the particles, namely,
@@
Here, $\mu=\frac{1}{2}(\alpha-\beta-\gamma+\alpha\beta-\beta\gamma)-\eta$. For further details, see Ref. Fiziev .
@@
Therefore, the harmonic oscillator with singular mass has quantized frequency.

### v2

@@ I Introduction
where $F(r,t)$ is the envelope function, $E(k)$ the dispersion energy, $k$ the crystal momentum, and $V(r)$ the potential due to external sources (impurities). Nonetheless, one highlights that considering a many-body Hamiltonian approximation to an electron, it is possible to derive a Schrödinger-like equation with PDM. In this context, the PDM concept has been the subject of intensive study in several systems with effective mass Mustafa , Costa , Zare , Ghafourian , Pourali . Furthermore, it is noteworthy that significant recent technological and theoretical advances have motivated considerable investigation into systems with PDM. Such interest arises from the crucial role played by this property, which enables the analysis of semiconductor impurities Slater , Luttinger , Wannier , nonlinear optical phenomena Guo , Zhang , and quantum-relativistic problems ILima0 . In addition to these applications, PDM problems are extensively used in studies concerning the influence of effective mass on polarons confined in parabolic wells RR1 , systems with complex optical potentials RR2 , as well as the quantum-mechanical description of systems subject to hyperbolic interactions RR3 , RR4 , among other RR5 , RR6 , RR8 , RR9 , RR10 , RR11 , RR12 . Although the study of problems involving PDM is a well-established field, we propose, for the first time, that a crystalline lattice, whose atoms inherently vibrate under the influence of a harmonic potential (see Ref. Kittel ), can exhibit singular characteristics when subjected to parallel stress induced by impurities, such as electromagnetic fields. This situation leads to the emergence of the singular mass profile, derived from multipolar contributions, as will be detailed in Eq. (8). In this framework, this work aims to investigate the quantum-mechanical description of the system resulting from the clustering of particles caused by such impurities. Although several studies have addressed problems related to singular PDM, we believe this is the first to explore the configuration of lattice defects using the PDM formalism.
@@ II The Generalized Schrödinger theory
Therefore, to avoid ambiguity arising from the symmetrization of the kinetic energy operator, allow us to consider Schrödinger’s equation exposed in Eq. (7). So, by analyzing Eq. (7) with the results presented in the literature, we conclude that our theory is equivalent to the ordered BenDaniel-Duke PDM problem BenDaniel . The BenDaniel-Duke theory has proven to be suitable for studying nanostructures subject to impurities, displacements, and geometric imperfections Willatzen . It is essential to highlight that our manuscript will focus on describing the quantum eigenstates; however, there are other works examining issues on the PDM Rev1 , Rev2 , Rev3 , Rev4 , Rev5 , Rev6 , Rev7 , Rev8 . For instance, one employs the PDM concept to investigate singular oscillator systems, viz., in which singularities arise due to singular interaction Rev1 . Furthermore, we use the PDM problems to describe several physical situations, e.g., transport of charge carriers in semiconductors with non-uniform composition Rev2 , and theoretical measures of quantum information Rev5 . Moreover, algebraic methods have shown great promise in the description of quantum eigenstates Rev3 .
@@ III The singular PDM
The singular mass emerges in various contexts, allowing for its interpretation from different perspectives. One exciting scenario for singular mass occurs in high-energy physics. In this framework, singular mass refers to a point in space where the gravitational field becomes infinitely intense, such as at the center of a black hole, where one believes that matter collapses into a singularity, resulting in infinite density Carroll , Maluf . Furthermore, the theory of singular mass also manifests in the fundamental principles of quantum mechanics Mustafa00 . In the quantum framework, singular mass is described through the motion of particles within the mass domain $m(x)$ Mustafa00 . Additionally, in quantum mechanics systems, a possible interpretation for PDM in quantum mechanical systems is regarded as a topological defect in the crystal lattice. Thus, one can understand the singular mass as a defect in the crystal lattice FLima .
In our study, we are interested in exploring the singular PDM physical implications of this system within a non-relativistic one-dimensional theory. Thus, let us consider a one-dimensional crystalline lattice as illustrated in Fig. 1.
@@
Here, $m_{0}$ has mass dimensions and describes the intensity of the effective mass. Furthermore, $x_{0}$ represents the equilibrium spacing between lattice sites. We present the PDM varying $m_{0}$ (keeping $x_{0}$ constant) and varying $x_{0}$ (keeping $m_{0}$ constant), respectively, in Fig. 2[(a) and (b)]. It is crucial to highlight that one can find some studies on the singular mass profiles in the literature Vakarchuk , ADAlhaidari , SMIkhdair . For instance, the singular PDM can describe an electron confined by electromagnetic fields. Indeed, in this framework, this mass profile arises due to the multipole expanse, see Ref. Vakarchuk .
@@ III.1 Connection with solid-state physics
Allow us to attempt to establish a phenomenological connection with our model. To accomplish our purpose, let us recall that the PDM has an equivalent distribution in reciprocal space. Thereby, the PDM at the reciprocal space  [^1] One defines the mass at reciprocal space as Fourier’s transform of the PDM at position space FLima . The calculation of mass in momentum space, within the context of solid-state physics, is significant, as this approach facilitates the understanding of the relationships between the electronic and mechanical properties of materials and their crystal structure Kittel . Furthermore, in momentum space, the dispersion energies of electrons are expressed as a function of the wave vector and the effective mass of charge carriers Kittel . is
@@ III.2 The eigenstates
Naturally, one assumes that the atoms in the crystalline lattice suffer the action of a restoring force Gomes , Kittel , i.e., the atoms belonging to the lattice behave as a set of harmonic oscillators. Consequently, a PDM at a crystalline lattice will submitted to a restoring force, i.e., a harmonic-like interaction. Hence, the potential will be
@@
Here, $\mu=\frac{1}{2}(\alpha-\beta-\gamma+\alpha\beta-\beta\gamma)-\eta$  [^2] It is pertinent to highlight that these conditions ensure that the eigenfunctions and their respective derivatives are continuous at the origin (i.e., $x=0$).. For further details, see Ref. Fiziev .
@@
Therefore, the harmonic oscillator with singular mass has quantized frequency. Here, one highlights that the quantization (or constraint) that arises in the frequency (33) is a consequence of the existence singularity in the mass distribution at $x=0$. Thus, this restriction is directly related to the energy’s real eigenvalues.